please help me with this problem!
A tugboat moves at a rate of 10 mph in still water. it travels 24 miles upstream and 24 miles downstream in a total time of 5 hours. what is the speed of the current?
Please and thank you.

Answer by jorel1380(3719)

24/(10+x)+24/(10-x)=5
24(10-x)+24(10+x)=5(100-x2)
240-24x+240+24x=500-5x2
5x2-20=0
x2-4=0
(x+2)(x-2)=0
x=-2 or 2
Throwing out the negative answer, we get the speed of the current to be 2mph..